OpenStudy (anonymous):

Find the slope of the line with the given points (6,0) and (4,8)

OpenStudy (anonymous):

still working on these ones eh?

OpenStudy (anonymous):

the formula y2-y1/x2-x1

OpenStudy (anonymous):

yea

myininaya (myininaya):

you have (x1=6,y1=0) and (x2=4, y2=8) just plug em in

OpenStudy (anonymous):

so, would y1 be 6, and y2 be 0?

OpenStudy (anonymous):

would the answer be m=-1.5?

myininaya (myininaya):

you could choose (x1=6,y1=0) and (x2=4, y2=8) or you can choose (x2=6,y2=0) and (x1=4,y1=8)

myininaya (myininaya):

no try again

myininaya (myininaya):

let me see you plug them into formula don't do anything else except plug them in

OpenStudy (anonymous):

8-0/4-6

myininaya (myininaya):

ok so after performing the operation on top you get=? and after performing the operation on bottom you get=?

OpenStudy (anonymous):

m=2?

myininaya (myininaya):

are you sure its just 2?

myininaya (myininaya):

is 4 greater than 6? yes so when you do 4-6 you get?

myininaya (myininaya):

oops the first question answer is no lol

myininaya (myininaya):

8-0=8 4-6=-2 so what is 8/(-2)=

OpenStudy (anonymous):

-2?

myininaya (myininaya):

-2*?=8

myininaya (myininaya):

-4?

OpenStudy (anonymous):

ok,-4

myininaya (myininaya):

-2*(-2)=4 not 8

OpenStudy (anonymous):

ok, so the answer is 4

myininaya (myininaya):

no

OpenStudy (anonymous):

@tc1979 u r great in division

OpenStudy (anonymous):

8/(-2)=-4

myininaya (myininaya):

right!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so the answer is -4

myininaya (myininaya):

yes you may need to work on your subtraction and division though if you want to get good at these

OpenStudy (anonymous):

so the slope is -4? that doesent make any sense

OpenStudy (anonymous):

usually slopes are in the form of a fraction

OpenStudy (anonymous):

use sense

myininaya (myininaya):

\[\frac{8-0}{4-6}=\frac{8}{-2}=-4\] -4 is a fraction -4/1

OpenStudy (anonymous):

well this one is -4/1 so it is still a fraction, but it is written in a different form. but you're right, most of the time it is a fraction

OpenStudy (anonymous):

so I can put -4 and it will be considered a fraction?

myininaya (myininaya):

you would write -4 since -4/1 is -4 lol

myininaya (myininaya):

i don;t know anymore this is getting too complex for me lol

OpenStudy (anonymous):

yep, if you write -4 it is a legitimate form of input

myininaya (myininaya):

jk